Mrs.Jonson just purchased a total of 12 books and magazines for her office. The total bill was $112. She spent $16 on magazines, and the books cost $12 each. How many books did she buy.
Answer by stanbon(75887)

Mrs.Jonson just purchased a total of 12 books and magazines for her office. The total bill was $112. She spent $16 on magazines, and the books cost $12 each. How many books did she buy.
Equation:
Value:: 12B + 16 = 112 dollars
----------------------------------
B = -4/3 + 28/3
----
B = 24/3 = 8 (# of books bought)
-----------
Cheers,
Stan H.
----------------